OpenStudy (anonymous):

Sammy is going to graph y = ex + 3. Before he graphs, he writes down what he thinks will happen to the parent graph. Which statement is correct

OpenStudy (anonymous):

We're missing some parts of this question. What are the multiple-choice answers?

OpenStudy (anonymous):

23 = 8 22 = 4 21 = 2 20 = 1 2-1 = 1 2 2-2 = 1 4 2-3 = 1 8 Look at the pattern and determine the value of x-2.

OpenStudy (anonymous):

A) x B) x2 C) 1 x2 D) 1 x